OpenStudy (anonymous):

PLEASE HELP!!! Jeannette's starting salary is $25,000 a year. She will get a 4% raise annually. How much money will she have made after 7 years? A. $165,824.39 B. $197,457.36 C. $230,355.66 D. $256,355.66

OpenStudy (john_es):

You have to sum the terms of a geometric series, \[a_n=25000\left(1+\frac{4}{100}\right)^{t-1}\] from t=1 to t=7. That gives, \[S=25000\frac{1-1.04^{t+1}}{1-1.04}=197457\]

OpenStudy (anonymous):

Thank you so much!!